OpenStudy (moazzam07):

Below are the data collected from two random samples of 100 members of a large travel club regarding the type of vacation they prefer: Sample Adventure Beach Cruise Ski A 22 73 2 3 B 20 71 5 4 Which of the following inferences can be made based on the data? Most members prefer a beach vacation. Most members prefer an adventure vacation. More members prefer a cruise vacation and a ski vacation than an adventure vacation. More members prefer an adventure vacation and a cruise vacation than a beach vacation.

OpenStudy (moazzam07):

Help please

OpenStudy (muscrat123):

is there a better way to show the data?

OpenStudy (moazzam07):

i dont think so

OpenStudy (muscrat123):

can u retype it? it is confusing bc the #s r everywhere

OpenStudy (moazzam07):

Sample Adventure Beach Cruise Ski A 22 73 2 3 B 20 71 5 4

OpenStudy (moazzam07):

the best i can do

OpenStudy (muscrat123):

what is the 22 and 20 for?

OpenStudy (moazzam07):

its part of the Graph

OpenStudy (muscrat123):

OH!!! ok. i think i might be able to help but mot sure yet.

OpenStudy (muscrat123):

ill try

OpenStudy (moazzam07):

ok thanks

OpenStudy (muscrat123):

for the answer do u pick one or more than one?

OpenStudy (moazzam07):

one

OpenStudy (muscrat123):

and u probably dont know how to solve right? thats y ur asking

OpenStudy (moazzam07):

yes

OpenStudy (muscrat123):

so first, u would start w/ the adventure.

OpenStudy (moazzam07):

ok o_o

OpenStudy (muscrat123):

i think u would need to find the average, 22 + 20 = 42 42/2=?

OpenStudy (moazzam07):

21

OpenStudy (moazzam07):

21

OpenStudy (muscrat123):

ok . so that is the average people who prefer adventure. do the same for beach

OpenStudy (moazzam07):

ok its 72 FOR beach

OpenStudy (moazzam07):

Cruise = 3.6

OpenStudy (moazzam07):

and ski is 3.5

OpenStudy (muscrat123):

cruise is wrong

OpenStudy (moazzam07):

3.5

OpenStudy (muscrat123):

yes!

OpenStudy (moazzam07):

2/5=0.4

OpenStudy (muscrat123):

2/5? what is that for?

OpenStudy (moazzam07):

you asked

OpenStudy (muscrat123):

when

OpenStudy (moazzam07):

the message you deletd

OpenStudy (moazzam07):

ok forget that ok what do i do next mate

OpenStudy (muscrat123):

u would find the greatest # of people i think ps:im not sure this is correct, but im pretty sure and also that is not what i asked. i deleted it bc u got it right after i typed it

OpenStudy (muscrat123):

ill draw what i meant

OpenStudy (moazzam07):

ok

OpenStudy (muscrat123):